Ten kilograms of steam at 100

0C is condensed in 500 kg of water

at 40 0C. What is the resulting temperature? (Ans: 51.8 0C


1
Expert's answer
2022-03-15T10:35:56-0400

Solution.

"m_s=10kg;"

"t_s=100^oC;"

"m_w=500kg;"

"t_w=40^oC;"

"r=2.3\sdot10^6J/kg;"

"c=4200 J\/{kg^oC};"

"t_f-?;"

"rm_s+cm_s(t_s-t_f)=cm_w(t_f-t_w);"

"rm_s+cm_st_s-cm_st_f=cm_wt_f-cm_wt_w;"

"rm_s+cm_st_s+cm_wt_w=cm_wt_f+cm_st_f;"

"t_f(cm_w+cm_s)=rm_s+cm_st_s+cm_wt_w;"

"t_f=\dfrac{rm_s+cm_st_s+cm_wt_w}{cm_w+cm_s};"

"t_f=\dfrac{2.3\sdot10^6\sdot10+4200\sdot10\sdot100+4200\sdot500\sdot40}{4200(10+500)}=51.9^oC;"

Answer: "t_f=51.9^oC."